OpenStudy (anonymous):

Two separate bacteria populations grow each month and are represented by the functions f(x) = 3x and g(x) = 7x + 6. In what month is the f(x) population greater than the g(x) population?

OpenStudy (anonymous):

@CBARREDO1

OpenStudy (anonymous):

MONTH 1 MONTH 2 MONTH 3 MONTH 4

OpenStudy (campbell_st):

just a quick question are the equations 1. f(x) = 3x g(x) = 7x + 6 or 2. \[f(x) = 3^x ~~~g(x) = 7x + 6\]

OpenStudy (anonymous):

2

OpenStudy (anonymous):

@campbell_st its the second equation

OpenStudy (anonymous):

month 2 is wrong

OpenStudy (campbell_st):

ok... that makes more sense... an easy method is to use a table of values  just keep going until f(x) is larger than g(x)